OpenStudy (anonymous):

simplify 3[ 25 - (-31 - 18)]

OpenStudy (zarkon):

222

OpenStudy (anonymous):

are you sure?

OpenStudy (zarkon):

http://alturl.com/x2w63

OpenStudy (zarkon):

yes ;)

OpenStudy (anonymous):

the way that zarkon achieved his answer is by doing the sums in the brackets first (-31-18)=-49 25-(-49)=25+49=74 3*74=222

OpenStudy (zarkon):

No I didn't...I used my calculator ;)